OpenStudy (anonymous):

PLEASE help me! I'm having trouble solving this raational inequality! x/(x+1)

OpenStudy (amistre64):

when are they equal?

OpenStudy (anonymous):

http://curriculum.kcdistancelearning.com/courses/ALG2x-HS-A09/a/assessments/Rt-SolvingRationalEquationsQuiz/A2_3.3_Quiz_24.gif I know I have to convert it into an equation x/(x+1) - x/(x-1) = 0 but then I have to multiply it by the LCM and I don't know how to calculate LCM for these values (x+1) and (x-1)

OpenStudy (amistre64):

forget the LCM, just cross multiply

OpenStudy (anonymous):

I get x(x-1) - x(x+1)

OpenStudy (amistre64):

good, now distribute thru

OpenStudy (anonymous):

2x^2 = 0?

OpenStudy (amistre64):

might need a little work on your distribution skills... x(x-1) - x(x+1) x^2-x - x^2-x -2x = 0

OpenStudy (anonymous):

The final answer is x=2 right?

OpenStudy (amistre64):

of course not: divide each side by -2 to get x= 0

OpenStudy (anonymous):

Ok. So the final answer is x>0?

OpenStudy (amistre64):

no, you might want to wait to finding the end, until we arrive at the end.

OpenStudy (anonymous):

Ok so x=0 how do we keep going? Substitute the values?

OpenStudy (amistre64):

lets draw an x axis ... mark it at 0 to define where they are equal; then run some simple tests

OpenStudy (amistre64):